OpenStudy (anonymous):

1) 26=8+v 2) 3+p=8 3)15+b=23

OpenStudy (anonymous):

I have 27 more step eqautions so HELP

Mehek (mehek14):

\(\tt{1)26-8=v\\2)8-3=9\\3)23-15=b}\) subtract for the answers

Mehek (mehek14):

\(\tt{2)8-3=p}\)*

OpenStudy (anonymous):

4) -15+n=-9 5) m+4=-12 6) x-7=13

Mehek (mehek14):

\(\tt{4)-9+15=n\\5)-12-4=m\\6)13+7=x}\)

OpenStudy (anonymous):

M-9= - 13

Mehek (mehek14):

\(\tt{-13+9}\)

OpenStudy (anonymous):

Can you hold on for about 3 minutes

Mehek (mehek14):

sure

OpenStudy (anonymous):

I need the most help on negatives

Mehek (mehek14):

ok

OpenStudy (anonymous):

Negatives have me a a 62 average in preap math so me =horable at negative eqautions

OpenStudy (anonymous):

Need help with number 4

Mehek (mehek14):

4) -15+n=-9?

OpenStudy (anonymous):

Yes that one

Mehek (mehek14):

so you would add \(\tt{-9+15}\) here's how u can remember subtract 15-9 = 6 which number is greater? 15 the sign for 15 is + so the answer is + +6

Mehek (mehek14):

or you can just switch the numbers \(\tt{-9+15\rightarrow15-9}\)

OpenStudy (anonymous):

15 is negative

OpenStudy (anonymous):

So -6

Mehek (mehek14):

when you move it to the other side, it changes the ign

Mehek (mehek14):

EX: -3 + x = 5 you want x by itself so you have to add 3 to cancel the 3 by the x

Mehek (mehek14):



OpenStudy (anonymous):

Ok sorry if this sounds stubborn i am no good at negitives so this is kind of getting me :/

Mehek (mehek14):



OpenStudy (anonymous):

Boh ok 5 m+4=-12

OpenStudy (anonymous):

12- 4 = 8

Mehek (mehek14):

-12 - 4

OpenStudy (anonymous):

- or +

OpenStudy (anonymous):

I think -

Mehek (mehek14):

-12 stays -12 +4 becomes -4 because it was moved to the other side -12 - 4 = -16

OpenStudy (anonymous):

You add it ?

Mehek (mehek14):

negative - negative = you can add but keep the answer negative

Mehek (mehek14):

like -2 - 2 = -4

OpenStudy (anonymous):

Ok so will you tutor me a little bit a day? I gtg

Mehek (mehek14):

same and yes I can

OpenStudy (anonymous):

Thx about 6pm?